# Equilibrium Portfolio Selection under Utility-Variance Analysis of Log Returns in Incomplete Markets

### v1

@@ Abstract
This paper investigates a time-inconsistent portfolio selection problem in the incomplete market model, integrating expected utility maximization with risk control. The objective functional balances the expected utility and variance on log returns, giving rise to time inconsistency and motivating the search of a time-consistent equilibrium strategy. We characterize the equilibrium via a coupled quadratic backward stochastic differential equation (BSDE) system and establish the existence theory in two special cases: (i) two Brownian motions driven the price dynamics and the factor process are independent with $\rho=0$; (ii) the trading strategy is constrained to be bounded. For the general case with correlation coefficient $\rho\neq 0$, we introduce the notion of an approximate time-consistent equilibrium. By employing the solution structure from the equilibrium in the case $\rho=0$, we can construct an approximate time-consistent equilibrium in the general case with an error of order $O(\rho^{2})$. Numerical examples and financial insights based on deep learning algorithms are also presented.
@@ 1 Introduction
The optimal portfolio management problem has always been a core research topic in quantitative finance, traditionally approached via the mean-variance (MV) criterion proposed by Markowitz_1952 and the expected utility (EU) theory developed in merton_1969; MERTON1971373. The classical mean–variance criterion captures the trade-off between the expected return and the risk (variance). In contrary, the utility maximization framework incorporates the investor’s risk aversion level into decision making without concerning the risk level of terminal wealth.
@@
Nevertheless, the log returns and the variance term in the objective functional render the problem time-inconsistent. That is, a strategy being optimal today may no longer remain optimal at future dates. Strotz1955 classifies agents facing time-inconsistency into three types. The naive agent always re-optimizes the current objective, ignoring time-inconsistency and thus constantly changing strategies. The precommitted agent fixes an initial optimal plan, knowing it will later become suboptimal; the plan remains time-inconsistent since it is only optimal at the initial time. A sophisticated agent cannot precommit but anticipates future deviations, viewing her future selves as strategic counterparts and seeking a time-consistent equilibrium strategy that each self optimally follows—namely, the intrapersonal equilibrium.
@@
Contrary to Wong_2017 and Bensouusan_2022, who focus on the precommitted strategy, the present paper aims to characterize a time-consistent equilibrium strategy by a sophisticated agent under game theoretical thinking with future selves, whose precise definition in continuous time was first proposed in Ekeland2006noncommitmentsubgameperfectequilibrium [^1] Since then, a strand of literature on time-inconsistent control problems in continuous-time setting has emerged, see e.g. Basak_dynamic_2010, hu_time-inconsistent_2012, hu_time-inconsistent_2017, Bjork_2014, dai_dynamic_2021 for MV problems, Ekeland2006noncommitmentsubgameperfectequilibrium, Ekeland_2007, Hamaguchi_2021 for non-exponential discounting problems.. Moreover, we consider the incomplete market model, where the financial market consists of a risk-free asset and a single risky asset [^2] For simplicity, we focus on one risky asset, as the analysis can be extended directly to multiple risky assets.. The wealth process is driven by a standard Brownian motion $B$, while the asset prices are influenced by an exogenous factor process driven by another Brownian motion $\bar{B}$. The correlation coefficient of these two Brownian motions is denoted by $\rho$. Similar to Hamaguchi_2021 and dai_dynamic_2021, we can employ the perturbation and maximum principle to establish a necessary condition to characterize an open-loop time-consistent equilibrium strategy in Theorem 3.1 for our utility maximization problem with variance minization via a two-dimensional system of coupled BSDEs (3.1) (see also (4.2) for a simplified yet still coupled system), which has not been addressed before in the existing literature.
@@
In general, the system (4.2) consists of two quadratic BSDEs coupled by the derived optimal strategy in quadratic form. Classical existence results for the multidimensional BSDE mainly fall into two categories: the fixed-point approach based on contraction mappings, which ensures uniqueness of solutions (see HU20161066, Luo_2017 and FAN2023105), and the construction of uniformly convergent approximations via coefficient regularization (see xing_class_2018, Jackson_2022 and JACKSON2023). However, neither of these methods applies to our problem (see Appendix C for the detailed explanation).
@@
For the general case with $\rho\neq 0$ and without trading constraint, we can no longer establish the existence of solution to the more complicated BSDE system. However, as a new contribution to the existing literature, we can show that by employing the solution structure from the time-consistent equilibrium strategy in the special case with $\rho=0$ we can successfully construct an approximate time-consistent equilibrium (see Definition 5.1) in the general case with $\rho\neq 0$ when the correlation coefficient $\rho$ is sufficiently small. More importantly, under some additional regularity conditions, we can rigorously show that the approximation error using the equilibrium solution in the case $\rho=0$ is of order $O(\rho^{2})$, see Theorem 5.5. This approximation result using the explicit equilibrium structure in the special case with $\rho=0$ significantly simplifies the decision making in some general incomplete market models with small $\rho$ such that it is sufficient to implement the more tractable control process while the time-consistency can be retained in an approximate sense.
@@
The remainder of the paper is organized as follows. Section 2 formulates the time inconsistent control problem as the trade-off between the utility maximization and the variance minimization over logarithmic portfolio returns, for which a time-consistent equilibrium is introduced. Section 3 provides the characterization of the time-consistent equilibrium by a quadratic BSDE system. Section 4 investigates the existence of solution to the BSDE system in the stochastic factor model in two special cases: (i) when two Brownian motions in the price dynamics and in the factor model are independent; (2) when the trading strategy is constrained in a compact set. Section 5 establishes the approximate time-consistent equilibrium in the general case when two Brownian motions are correlated with $\rho\neq 0$ using the same solution structure from time-consistent equilibrium solution in the model with $\rho=0$ with the error of order $O(\rho^{2})$. Some numerical illustrations and financial implications are presented in Section 6. Finally, Appendix A and B collect some inequalities and sufficient conditions for the existence theory of BSDE system that are used in previous sections, and Appendix C elaborates some challenges of the BSDE system in the general case when $\rho\neq 0$.
Notations: For $t\in[0,T]$, $p,q\geq 1$, we list below some notations that will be used frequently throughout the paper:
@@
    $\mathcal{T}_{[0,T]}$ denotes all stopping times on $[0,T]$.
  • $L^{\infty}_{\mathcal{F}_{t}}(\Omega;\mathbb{R})$ denotes the set of all $\mathbb{R}$-valued $\mathcal{F}_{t}$-measurable bounded random variables.
  • $L^{p}_{\mathbb{F}}(\Omega;L^{q}(t,T;\mathbb{R}))$ denotes the set of all $\mathbb{R}$-valued progressively measurable processes $X$ with
@@
    When $p=q$, we write $L^{p}_{\mathbb{F}}(t,T;\mathbb{R}):=L^{p}_{\mathbb{F}}(\Omega;L^{p}(t,T;\mathbb{R}))$
  • $L^{p}_{\mathbb{F}}(\Omega;C([t,T];\mathbb{R}))$ denotes the set of all $\mathbb{R}$-valued adapted and continuous processes $X$ with
@@
  • $\mathcal{S}^{\infty}$ denotes the set of all $\mathbb{R}$-valued $\mathbb{F}$-progressively measurable and bounded processes.
@@ 2 Problem Formulation
Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t},0\leq t\leq T\}$ is the augmented natural filtration generated by two standard Brownian motions $\{B(t),0\leq t\leq T\}$ and $\{\bar{B}(t),0\leq t\leq T\}$. Their correlation coefficient $\rho$ is such that $\mathbb{E}[B(t)\bar{B}(t)]=\rho t$. The market model consists of one risk-free bond with interest rate $r=\{r_{t}:t\in[0,T]\}$ and one stockwith price dynamics governed by
@@
where the drift $\mu:[0,T]\times\Omega\to\mathbb{R}$, the volatility $\mathbf{\sigma}>0:[0,T]\times\Omega\to\mathbb{R}$, and the interest rate $r:[0,T]\times\Omega\to\mathbb{R}$, are all $\mathbb{F}$-progressively measurable processes.
@@
$r_{t},\mu_{t},\sigma_{t}>0$ and $\theta_{t}:=\sigma_{t}^{-1}(\mu_{t}-r_{t})$ are bounded $\mathbb{F}$-progressively measurable processes.
A trading strategy is a $\mathbb{F}$-progressively measurable processes $\pi$ such that $\int_{0}^{T}|\pi_{t}(\mu(t)-r(t))|\mathrm{d}t+\int_{0}^{T}|\sigma(t)\pi_{t}|^{2}\mathrm{d}t<\infty$ a.s., where $\pi_{t}$ stands for the fraction of wealth allocated in the stock at time $t$. The self-financing wealth process $W^{\pi}$ evolves according to
@@
In this paper, we consider a utility maximization framework with variance adjustment on log return of the portfolio. That is, with given $r_{0}\in\mathbb{R},\pi_{0}$ and $t\in[0,T)$, the objective functional of the investor given the trading strategy $\pi$ at time $t$ is defined by
@@
where $\mathbb{E}_{t}$ and $\textup{Var}_{t}$ represent the conditional expectation and the conditional variance at time $t$, respectively, and $\gamma_{t}>0$ can be regarded as the weighting parameter measuring the trade-off between the expected utility and the risk (variance) at time $t$. In particular, we shall work with the CARA utility $U(x)=-\frac{1}{\zeta}e^{-\zeta x}$ with the risk aversion parameter $\zeta>0$.
To simplify the notation, let us denote $R^{\pi}_{t}=\log W^{\pi}_{t}$, which satisfies
$\mathrm{d}R^{\pi}_{t}=\left[r_{t}+\pi_{t}(\mu_{t}-r_{t})-\frac{1}{2}|\sigma_{t}\pi_{t}|^{2}\right]\mathrm{d}t+\pi_{t}\sigma_{t}\mathrm{d}B_{t},\quad R^{\pi}_{0}=r_{0}=\log w_{0}\in\mathbb{R}.$
Then the objective functional in (2.1) at time $t$ can be written as
$\mathbb{E}_{t}\left[U(R^{\pi}_{T}-R^{\pi}_{t})\right]-\frac{\gamma_{t}}{2}\textup{Var}_{t}[R^{\pi}_{T}-R^{\pi}_{t}]=-\frac{1}{\zeta}\mathbb{E}_{t}\left[\exp\left(-\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]-\frac{\gamma_{t}}{2}\textup{Var}_{t}[R^{\pi}_{T}].$
To ensure the objective functional under $\pi$ is well defined, we impose the following condition on trading strategies:
@@
(Admissible trading strategy) A trading strategy $\pi$ is admissible if $\pi\in H_{\textup{BMO}}$ and, for some $p>1$, it satisfies
$\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]<\infty,\quad a.s.$ (2.2)
for any $t\in[0,T)$. We denote by $\Pi_{0}$ the set of all admissible trading strategies.
@@
Note that we have to require $\mathbb{E}_{t}\left[\exp\left(-\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]<\infty,a.s.$ to ensure that the objetive functional at time $t$ is well-defined. In (2.2), we impose a slightly stronger integrability condition (with $p>1$). As will be shown in the subsequent discussion, this condition guarantees that the admissible set remains open under perturbations (see Lemma 2.7 and Corollary 2.9), which is essential for the rigorous definition of the equilibrium. Moreover, it also implies that $\mathbb{E}_{t}\left[\left(U^{\prime}(R^{\pi}_{T}-R^{\pi}_{t})\right)^{p}\right]<\infty,a.s.$, a fact that will play a crucial role in the proof of Theorem 3.1; see Hamaguchi_2021 for similar conditions.
The following lemma shows that $J(t,\pi)<\infty$ for any $\pi\in\Pi_{0}$.
@@
For $\pi\in\Pi_{0}$ and $t\in[0,T)$, $J(t,\pi)$ is finite almost surely. Moreover, $R^{\pi}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$.
@@
First, by Hölder’s inequality, we have
$\mathbb{E}_{t}\left[\exp\left(-\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]\leq\left(\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]\right)^{\frac{1}{p}}<\infty,\quad a.s..$
By the definition of the admissible strategy $\pi$, it holds that
Using the standard estimate for SDE (see Theorem 3.4.3 in zhang_backward_2017), we obtain that
$\mathbb{E}\left[\sup\limits_{s\in[t,T]}\left|R^{\pi}_{s}-R^{\pi}_{t}\right|^{2}\mathds{1}_{A}\right]=$ $C(T)\mathbb{E}\left[\left(\int_{t}^{T}\left[r_{s}+\pi_{s}(\mu_{s}-r_{s})-\frac{1}{2}|\sigma_{s}\pi_{s}|^{2}\right]\mathds{1}_{A}\mathrm{d}s\right)^{2}+\int_{t}^{T}|\pi_{s}\sigma_{s}|^{2}\mathds{1}_{A}\mathrm{d}s\right]$
$\leq$ $C(\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T)\mathbb{E}\left[\left(\left(\int_{t}^{T}|\pi_{s}|^{2}\mathrm{d}s\right)^{2}+1\right)\mathds{1}_{A}\right].$
Note that $A\in\mathcal{F}_{t}$ is arbitrary and $\pi\in H_{\textup{BMO}}$, by the definition of conditional expectation and Lemma A.1, we deduce that
$\left(\mathbb{E}_{t}\left[\left|R^{\pi}_{T}-R^{\pi}_{t}\right|\right]\right)^{2}\leq\mathbb{E}_{t}\left[\left|R^{\pi}_{T}-R^{\pi}_{t}\right|^{2}\right]$ $\leq\mathbb{E}_{t}\left[\sup\limits_{s\in[t,T]}\left|R^{\pi}_{s}-R^{\pi}_{t}\right|^{2}\right]$
$\leq C(\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T)\left(\mathbb{E}_{t}\left[\left(\int_{t}^{T}|\pi_{s}|^{2}\mathrm{d}s\right)^{2}\right]+1\right)$
$\leq C(\|r\|_{\infty},\|\mu\|_{\infty},\|\sigma\|_{\infty},T)\left(\|\pi\|_{\textup{BMO}}^{4}+1\right)<\infty.$
Thus $\textup{Var}_{t}[R^{\pi}_{T}-R^{\pi}_{t}]$ is bounded, and consequently, $J(t,\pi)$ is finite for any $t\in[0,T)$. The proof also verifies that $R^{\pi}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$. ∎
Remark 2.5.
Note that for a given admissible trading strategy $\pi$, $R^{\pi}_{T}-R^{\pi}_{t}$ is independent of the initial value $r_{0}$. Hence, the definition of the admissible set $\Pi_{0}$ is independent of the initial value $r_{0}$.
The issue of time inconsistency arises due to the initial dependence of $R_{t}$ and the variance term, as a result, the optimal solution make no longer sense as the optimal decision today might be turned over at future dates. Instead, we seek the equilibrium solution using the game theoretic thinking with future selves. To define the equilibrium strategy, we first introduce the perturbation strategy of an admissible strategy.
@@
Assume that $\pi$ is an admissible strategy, for any fixed $t\in[0,T)$, $\varepsilon\in(0,T-t)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$, define an admissible perturbation strategy $\pi^{t,\varepsilon,\eta}$ (see Corollary 2.9) as follows:
@@
Let $\xi^{t,\varepsilon,\eta,\pi}:=R^{\pi^{t.\varepsilon,\eta}}-R^{\pi}$ be the difference of log return processes between $\pi^{t,\varepsilon,\eta}$ and $\pi$. Then for any $t\in[0,T)$, $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R}^{d})$, $\varepsilon\in(0,T-t)$ and an admissible strategy $\pi$, $\xi^{t,\varepsilon,\pi}:=\xi^{t,\varepsilon,\eta,\pi}$ [^3] For notation simplicity, we drop the superscript $\eta$. is the unique solution of the SDE [^4] Formally, both $R^{\pi^{t.\varepsilon,\eta}}$ and $\xi^{t,\varepsilon,\pi}$ can be defined on the entire interval $[0,T]$. Nevertheless, in view of the definition of equilibrium, it suffices to consider them only on $[t,T]$. that
$\left\{\begin{aligned} &\mathrm{d}\xi_{s}^{t,\varepsilon,\pi}=\left(a(s,\pi^{t,\varepsilon,\eta}_{s})-a(s,\pi_{s})\right)\mathrm{d}s+\sigma_{s}\eta 1_{s\in[t,t+\varepsilon)}\mathrm{d}B_{s},\\ &\xi_{t}^{t,\varepsilon,\pi}=0.\end{aligned}\right$. (2.3)
Here, $a(s,\pi):=r_{s}+(\mu_{s}-r_{s})\pi-\frac{1}{2}|\sigma_{s}\pi|^{2}$ . In the following lemma, we will first establish some fundamental properties of $\xi^{t,\varepsilon,\pi}$.
@@
Let $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$ be fixed.

### v2

@@ Abstract
This paper investigates a time-inconsistent portfolio selection problem in the incomplete market model, integrating expected utility maximization with risk control. The objective functional balances the expected utility and variance on log returns, giving rise to time inconsistency and motivating the search for a time-consistent equilibrium strategy. We characterize the equilibrium via a coupled quadratic backward stochastic differential equation (BSDE) system and establish the existence results in two special cases: (i) the two Brownian motions driving the price dynamics and the factor process are independent ($\rho=0$); (ii) the trading strategy is constrained to be bounded. For the general case with correlation coefficient $\rho\neq 0$, we introduce the notion of an approximate time-consistent equilibrium. Employing the solution structure from the equilibrium in the case $\rho=0$, we can construct an approximate time-consistent equilibrium in the general case with an error of order $O(\rho^{2})$. Numerical examples and financial insights are also presented based on deep learning algorithms.
@@ 1 Introduction
The optimal portfolio management problem has always been a core research topic in quantitative finance, traditionally approached via the mean-variance (MV) criterion proposed by Markowitz_1952 and the expected utility (EU) theory developed in merton_1969; MERTON1971373. The classical MV criterion captures the trade-off between the expected return and the risk (variance). In contrary, the utility maximization framework incorporates the investor’s risk aversion level into decision making without concerning the risk level of terminal wealth.
@@
Nevertheless, the log returns and the variance term in the objective functional render the problem time-inconsistent. That is, a strategy being optimal today may no longer remain optimal at future dates. Strotz1955 classifies agents facing time-inconsistency into three types. The naive agent always re-optimizes the current objective, ignoring time-inconsistency and thus constantly changing strategies. The precommitted agent fixes an initial optimal plan, knowing it will later become suboptimal; the plan remains time-inconsistent because it is only optimal at the initial time. A sophisticated agent cannot precommit but anticipates future deviations, viewing her future selves as strategic counterparts and seeking a time-consistent equilibrium strategy that each self optimally follows—namely, the intrapersonal equilibrium.
@@
Contrary to Wong_2017 and Bensouusan_2022, who focus on the precommitted strategy, the present paper aims to characterize a time-consistent equilibrium strategy by a sophisticated agent under game theoretical thinking with future selves, whose precise definition in continuous time was first proposed in Ekeland2006noncommitmentsubgameperfectequilibrium [^1] Since then, a strand of literature on time-inconsistent control problems in continuous-time setting has emerged, see e.g. Basak_dynamic_2010, hu_time-inconsistent_2012; hu_time-inconsistent_2017, Bjork_2014, dai_dynamic_2021 for MV problems, Ekeland2006noncommitmentsubgameperfectequilibrium, Ekeland_2007, Hamaguchi_2021 for non-exponential discounting problems.. Moreover, we consider the incomplete market model, where the financial market consists of a risk-free asset and a single risky asset [^2] For simplicity, we focus on one risky asset, as the analysis can be extended directly to multiple risky assets.. The wealth process is driven by a standard Brownian motion $B$, while the asset prices are influenced by an exogenous factor process driven by another Brownian motion $\bar{B}$. The correlation coefficient of these two Brownian motions is denoted by $\rho$. Similar to Hamaguchi_2021 and dai_dynamic_2021, we can employ the perturbation and calculus of variation method to establish a necessary condition to characterize an open-loop time-consistent equilibrium strategy in Theorem 3.1 for our utility maximization problem with variance minization via a two-dimensional system of coupled BSDEs (3.1) (see also (4.2) for a simplified yet still coupled system), which has not been addressed before in the existing literature.
@@
In general, the system (4.2) consists of two quadratic BSDEs coupled by the derived equilibrium strategy in quadratic form. Classical existence results for the multidimensional quadratic BSDEs mainly fall into two categories: the fixed-point approach based on contraction mappings, which ensures uniqueness of solutions (see HU20161066, Luo_2017 and FAN2023105), and the construction of uniformly convergent approximations via coefficient regularization (see xing_class_2018, Jackson_2022 and JACKSON2023). However, neither of these methods applies to our problem (see Appendix C for the detailed explanation).
@@
For the general case with $\rho\neq 0$ and without trading constraint, we can no longer establish the existence of solution to the more complicated BSDE system. However, as a new contribution to the existing literature, we show that by employing the solution structure from the time-consistent equilibrium strategy in the special case with $\rho=0$, one can successfully construct an approximate time-consistent equilibrium (see Definition 5.1) in the general case with $\rho\neq 0$ when the correlation coefficient $\rho$ is sufficiently small. More importantly, under some additional regularity conditions, we can rigorously show that the approximation error using the approximate time-consistent equilibrium of order $O(\rho^{2})$, see Theorem 5.6. This approximation result using the explicit equilibrium structure in the special case with $\rho=0$ significantly simplifies the decision making in some general incomplete market models with small $\rho$ such that it is sufficient to implement the more tractable control process while the time-consistency can be retained in an approximate sense.
@@
The remainder of the paper is organized as follows. Section 2 formulates the time inconsistent control problem as the trade-off between the utility maximization and the variance minimization over logarithmic portfolio returns, for which a time-consistent equilibrium is introduced. Section 3 provides the characterization of the time-consistent equilibrium by a quadratic BSDE system. Section 4 investigates the existence of solution to the BSDE system in the stochastic factor model in two special cases: (i) when two Brownian motions in the price dynamics and in the factor model are independent; (ii) when the trading strategy is constrained in a compact set. Section 5 establishes the approximate time-consistent equilibrium in the general case when two Brownian motions are correlated with $\rho\neq 0$ using the same solution structure from time-consistent equilibrium solution in the model with $\rho=0$ with the error of order $O(\rho^{2})$. Some numerical illustrations and financial implications are presented in Section 6. Section 7 concludes the paper with some remarks. Appendix A and B collect some inequalities and sufficient conditions for the existence theory of BSDE system that are used in previous sections, Appendix C elaborates some challenges of the BSDE system in the general case when $\rho\neq 0$, and Appendix D collects the proofs of some auxiliary results in previous sections.
Notations: Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t},0\leq t\leq T\}$ is the augmented natural filtration generated by two standard Brownian motions $\{B(t),0\leq t\leq T\}$ and $\{\bar{B}(t),0\leq t\leq T\}$. Their correlation coefficient $\rho$ is such that $\mathbb{E}[B(t)\bar{B}(t)]=\rho t$. For $t\in[0,T]$, $p,q\geq 1$, we list below several notations that will be used frequently throughout the paper:
@@
    and $\mathcal{T}_{[0,T]}$ denotes the set of all $\mathbb{F}$-stopping times taking values in $[0,T]$.
  • $L^{\infty}_{\mathcal{F}_{t}}(\Omega;\mathbb{R})$ denotes the set of all $\mathbb{R}$-valued, $\mathcal{F}_{t}$-measurable, bounded random variables.
  • $L^{p}_{\mathbb{F}}(\Omega;L^{q}(t,T;\mathbb{R}))$ denotes the set of all $\mathbb{R}$-valued, $\mathbb{F}$-progressively measurable processes $X$ such that
@@
    When $p=q$, we write $L^{p}_{\mathbb{F}}(t,T;\mathbb{R}):=L^{p}_{\mathbb{F}}(\Omega;L^{p}(t,T;\mathbb{R}))$.
  • $L^{p}_{\mathbb{F}}(\Omega;C([t,T];\mathbb{R}))$ denotes the set of all $\mathbb{R}$-valued, $\mathbb{F}$-adapted, continuous processes $X$ such that
@@
  • $\mathcal{S}^{\infty}$ denotes the set of all $\mathbb{R}$-valued, $\mathbb{F}$-progressively measurable, bounded processes.
  • A function $u$ is said to belong to $C^{2+\beta}(\mathbb{R}^{n})$ if and only if
    $\sup_{x\in\mathbb{R}^{n}}|u|+\sup_{x\in\mathbb{R}^{n}}|u_{x}|+\sup_{x\in\mathbb{R}^{n}}|u_{xx}|+\sup_{x,x^{\prime}\in\mathbb{R}^{n}}\frac{|u_{xx}(x)-u_{xx}(x^{\prime})|}{|x-x^{\prime}|^{\beta}}<\infty,$
    and a function $u$ is said to belong to $\in C^{1+\beta/2,\,2+\beta}([0,T]\times\mathbb{R}^{n})$ if and only if
    $\sup_{(t,x)\in[0,T]\times\mathbb{R}^{n}}|u|+\sup_{(t,x)\in[0,T]\times\mathbb{R}^{n}}|u_{t}|+\sup_{(t,x)\in[0,T]\times\mathbb{R}^{n}}|u_{x}|+\sup_{(t,x)\in[0,T]\times\mathbb{R}^{n}}|u_{xx}|$
    $+\sup_{(t,x),(t^{\prime},x^{\prime})\in[0,T]\times\mathbb{R}^{n}}\frac{|u_{xx}(t,x)-u_{xx}(t^{\prime},x^{\prime})|}{d\big((t,x),(t^{\prime},x^{\prime})\big)^{\beta}}+\sup_{(t,x),(t^{\prime},x^{\prime})\in[0,T]\times\mathbb{R}^{n}}\frac{|u_{t}(t,x)-u_{t}(t^{\prime},x^{\prime})|}{d\big((t,x),(t^{\prime},x^{\prime})\big)^{\beta}}<\infty,$
    where $d\big((t,x),(t^{\prime},x^{\prime})\big):=|x-x^{\prime}|+|t-t^{\prime}|^{1/2}$ is the parabolic distance between $(t,x)$ and $(t^{\prime},x^{\prime})$.
@@ 2 Problem Formulation
We consider a financial market consisting of one risk-free bond with interest rate $r=\{r_{t}:t\in[0,T]\}$ and one risky asset whose price dynamics are given by
@@
where the drift $\mu$, the volatility $\sigma$, and the interest rate $r$ are all $\mathbb{F}$-progressively measurable processes.
Throughout the paper, we impose the following assumption on the market coefficients.
@@
The processes $r,\mu$ and $\sigma$ are bounded. Moreover, $\mu>r>0,\sigma>0$ for almost every $t\in[0,T]$, $\mathbb{P}$-a.s., and $\sigma^{-1}$ is also bounded.
Under Assumption 2.1, the market price of risk $\theta:=\sigma^{-1}(\mu-r)>0$ are bounded.
A trading strategy is a $\mathbb{F}$-progressively measurable processes $\pi$ such that $\int_{0}^{T}|\pi_{t}(\mu(t)-r(t))|\mathrm{d}t+\int_{0}^{T}|\sigma(t)\pi_{t}|^{2}\mathrm{d}t<\infty$ a.s., where $\pi_{t}$ represents the proportion of wealth invested in the risky asset at time $t$. Given such a strategy, the corresponding self-financing wealth process $W^{\pi}$ evolves according to
@@
[…]
@@
where $\mathbb{E}_{t}$ and $\textup{Var}_{t}$ represent the conditional expectation and the conditional variance at time $t$, respectively, and $\gamma_{t}>0$ is the weighting parameter capturing the trade-off between expected utility and risk (variance) at time $t$. In particular, we shall work with the CARA utility $U(x)=-\frac{1}{\zeta}e^{-\zeta x}$ with the risk aversion parameter $\zeta>0$.
To simplify the notation, let us denote $R^{\pi}_{t}:=\log W^{\pi}_{t}$ as the log-return process, which satisfies
$\mathrm{d}R^{\pi}_{t}=\left[r_{t}+\pi_{t}(\mu_{t}-r_{t})-\frac{1}{2}|\sigma_{t}\pi_{t}|^{2}\right]\mathrm{d}t+\pi_{t}\sigma_{t}\mathrm{d}B_{t},\quad R^{\pi}_{0}=r_{0}=\log w_{0}\in\mathbb{R}.$ (2.2)
Under this notation, the objective functional (2.1) can be rewritten as
$J(t,\pi)=\mathbb{E}_{t}\left[U(R^{\pi}_{T}-R^{\pi}_{t})\right]-\frac{\gamma_{t}}{2}\textup{Var}_{t}[R^{\pi}_{T}-R^{\pi}_{t}]=-\frac{1}{\zeta}\mathbb{E}_{t}\left[\exp\left(-\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]-\frac{\gamma_{t}}{2}\textup{Var}_{t}[R^{\pi}_{T}].$
To ensure that the objective functional is well defined for the trading strategy $\pi$, we introduce the following admissibility condition.
@@
(Admissible trading strategy) For any $p>1$, define
$\Pi_{p}:=\left\{\pi\in H_{\textup{BMO}}:\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]<\infty\,\,a.s.\text{ for all}\,\,t\in[0,T)\right\}.$
A trading strategy is said to be admissible if it belongs to $\Pi_{p}$ for some $p>1$. And denote by $\Pi:=\bigcup\limits_{p>1}\Pi_{p}$ the collection of all admissible trading strategies.
@@
Note that for any trading strategy $\pi$, $R^{\pi}_{T}-R^{\pi}_{t}$ is independent of the initial value $r_{0}$. Thus, the set $\Pi$ does not depend on $r_{0}$. Furthermore, we require $\mathbb{E}_{t}\left[\exp\left(-\zeta(R^{\pi}_{T}-R^{\pi}_{t})\right)\right]<\infty$ a.s. to ensure that the objective functional at time $t$ is well-defined. In Definition 2.2, we impose a slightly stronger integrability requirement with a power $p>1$. As will be shown later, this strengthened condition guarantees that $\Pi$ remains stable under perturbations (see Lemma 2.6 and Corollary 2.8), which is crucial for a rigorous formulation of equilibrium strategies. Moreover, it implies that $\mathbb{E}_{t}\left[\left(U^{\prime}(R^{\pi}_{T}-R^{\pi}_{t})\right)^{p}\right]<\infty$ a.s, a property that plays an essential role in the proof of Theorem 3.1; see. e.g., Hamaguchi_2021 for similar integrability conditions.
The following lemma shows $J(t,\pi)<\infty$ a.s. for any $\pi\in\Pi$.
@@
For any $\pi\in\Pi$ and $t\in[0,T)$, $J(t,\pi)<\infty$ a.s.. Moreover, $R^{\pi}\in L^{2}_{\mathbb{F}}(\Omega;C([0,T];\mathbb{R}))$.
@@
See Appendix D.1. ∎
The issue of time inconsistency arises due to the initial dependence of $R_{t}$ and the presence of the variance term. As a result, the classical optimal solution may no longer be meaningful because a decision that appears optimal today may be overturned at future dates. Instead, we seek an equilibrium solution using a game-theoretic approach with future selves.
To define the equilibrium strategy, we first introduce the notion of a perturbed strategy for $\pi\in\Pi$.
@@
For any $\pi\in\Pi$, $t\in[0,T)$, $\varepsilon\in(0,T-t)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$, define the perturbed strategy $\pi^{t,\varepsilon,\eta}\in\Pi$ (see Corollary 2.8 for the proof of admissibility) by
@@
For any $t\in[0,T)$, $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R}^{d})$, $\varepsilon\in(0,T-t)$ and $\pi\in\Pi$, let $\xi^{t,\varepsilon,\eta,\pi}:=R^{\pi^{t.\varepsilon,\eta}}-R^{\pi}$ be the difference of log return processes between $\pi^{t,\varepsilon,\eta}$ and $\pi$. Then, $\xi^{t,\varepsilon,\pi}:=\xi^{t,\varepsilon,\eta,\pi}$ [^4] For notation simplicity, we drop the superscript $\eta$. is the unique solution to the SDE [^5] Formally, both $R^{\pi^{t.\varepsilon,\eta}}$ and $\xi^{t,\varepsilon,\pi}$ can be defined on the entire interval $[0,T]$. Nevertheless, in view of the definition of equilibrium, it suffices to consider them only on $[t,T]$.
$\left\{\begin{aligned} &\mathrm{d}\xi_{s}^{t,\varepsilon,\pi}=\left(a(s,\pi^{t,\varepsilon,\eta}_{s})-a(s,\pi_{s})\right)\mathrm{d}s+\sigma_{s}\eta 1_{s\in[t,t+\varepsilon)}\mathrm{d}B_{s},\quad s\in[t,T],\\ &\xi_{t}^{t,\varepsilon,\pi}=0,\end{aligned}\right$. (2.3)
where $a(s,\pi):=r_{s}+(\mu_{s}-r_{s})\pi-\frac{1}{2}|\sigma_{s}\pi|^{2}$ . In the following lemma, we establish several fundamental properties of $\xi^{t,\varepsilon,\pi}$.
@@
Let $\pi\in\Pi$, $t\in[0,T)$ and $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R})$ be fixed. Then we have the following three assertions: